# Geometry-Guided Ray Augmentation for Neural Surface Reconstruction with Sparse Views

### v1

@@
Abstract
In this paper, we propose a novel method for 3D scene and object reconstruction from sparse multi-view images. Different from previous methods that leverage extra information such as depth or generalizable features across scenes, our approach leverages the scene properties embedded in the multi-view inputs to create precise pseudo-labels for optimization without any prior training. Specifically, we introduce a geometry-guided approach that improves surface reconstruction accuracy from sparse views by leveraging spherical harmonics to predict the novel radiance while holistically considering all color observations for a point in the scene. Also, our pipeline exploits proxy geometry and correctly handles the occlusion in generating the pseudo-labels of radiance, which previous image-warping methods fail to avoid. Our method, dubbed Ray Augmentation (RayAug), achieves superior results on DTU and Blender datasets without requiring prior training, demonstrating its effectiveness in addressing the problem of sparse view reconstruction. Our pipeline is flexible and can be integrated into other implicit neural reconstruction methods for sparse views.
1 Introduction
Figure 1: With a limited number of posed images (left), our proposed RayAug could produce high-quality novel view synthesis and accurate surface reconstruction when training on clean images (right). We reconstruct the meshes using the marching cubes algorithm [17].
Reconstructing 3D geometry from multi-view inputs is a fundamental and challenging problem in computer graphics and computer vision. Traditional multi-view stereo (MVS) methods primarily rely on correspondence matching between input images, followed by fusing predicted depth maps into point clouds for the final meshing process. However, these techniques require hand-crafted heuristics, which may struggle to find accurate matches that are crucial for producing meaningful dense surfaces when dealing with sparse input images.
Recent advances in neural implicit representations have shown promising results for 3D scene representation. The groundbreaking neural radiance fields (NeRF) [19] leverages a multi-layer perceptron (MLP) to predict per-point radiance and density and uses differentiable volume rendering to synthesize novel views. Neural surface reconstruction [32, 35, 36] methods also leverage radiance fields based techniques to supervise view synthesis for surface reconstruction. However, like traditional approaches, they also require dense images to reconstruct intricate geometry, which limits their applicability.
To address the sparse view problem, some methods have been proposed that enhance supervision during training. These approaches primarily use multi-view consistency to create pseudo labels for unseen rays by re-projection [39] or image warping [33] 3D points to unseen viewpoints. While these techniques improve results in sparse view settings, their augmentation strategies face two key issues: (1) they do not guarantee that the reprojected or warped points are visible at the augmented camera poses, and (2) they simply assign the color of the augmented rays based on the source view, leading to inaccurate estimations.
In this paper, we address these limitations and develop a method to generate more precise pseudo-labels for training radiance fields. We empirically observe that the rough geometry in the early stage of the implicit surface representation training provides valuable information with unbiased surface estimation and regularization. By combining this insight with physical priors, we propose RayAug, a novel geometry-guided approach for neural reconstruction from sparse views.
Specifically, we shoot rays from 3D points on the surface of the rough geometry to unseen viewpoints, ensuring their observability by ray marching the signed-distance field. Rays that hit the surface are discarded. Additionally, since a 3D point can be seen from multiple input viewpoints, we aggregate all the color observations by fitting spherical harmonics (SH) for each spatial point. This approach aligns with the distribution of natural radiance, providing better extrapolation results than linear methods. The color of each augmented ray from the same 3D point is then obtained by querying the fitted spherical harmonics function for each point. Finally, we perform another visibility check to avoid biased estimation of color caused by occlusion. We evaluate our method through novel view synthesis and mesh reconstruction on DTU dataset [10] and Blender dataset [19], demonstrating significant improvements compared to baseline methods.
In summary, our contributions are as follows:
  • We exploit spherical harmonics to holistically consider all color observations for a 3D point, enabling pseudo-label supervisions consistent with the natural distribution of radiance.
  • We propose to use rough geometry to avoid occlusion in the ray augmentation process, which avoids the confound of point radiance and reconstructs more accurate surfaces from sparse views.
@@ 1 Introduction
  • Our method works as a plug-and-play method and can be incorporated with other techniques for sparse implicit neural surface reconstruction. Our method achieves superior results on sparse view reconstruction without the need for extra data and prior training.
2 Related Work
2.1 Multi-view Reconstruction
The traditional multi-view stereo (MVS) pipeline first estimates per-view depth maps by incorporating and matching images. Then the depth maps are filtered and fused to obtain a global representation i.e. point clouds [7, 27]. 3D surfaces can thus be reconstructed by applying meshing algorithms such as Poisson reconstruction [13] or Delaunay triangulation [15] on point clouds. Later, learning-based approaches provide more accurate depth estimation by exploiting massive data and achieving better performance on MVS tasks [34, 8]. These methods produce promising results with dense captured images [26], but rely heavily on the quality of correspondence matching. Therefore, they are sensitive to outliers, noises, and textureless regions, leading to incomplete and noisy reconstructions.
2.2 Neural Surface Reconstruction
@@
Recently, neural implicit representations have demonstrated superior effectiveness in various tasks such as 3D object [23, 18, 3] and scene representation [19, 29, 28], novel view synthesis [19, 1, 31] and multi-view 3D reconstruction [35, 21, 32, 22]. Among them, NeRF and its variants [19, 1] demonstrated state-of-the-art performance for novel view synthesis with differentiable volumetric rendering. However, they require dense images for input and it is hard to extract high-quality geometry because the predicted density field lacks sufficient constraints. Other works utilizing surface rendering [21] produce more accurate modeling but require additional constraints such as ground truth masks or depth priors for supervision [38, 35]. Several later studies combine the advantages of surface and volume representation, generating satisfying surfaces with only image inputs [32, 36, 22]. For example, NeuS [32] combined volume rendering with signed distance functions to circumvent the problem of insufficient surface constraint in the original NeRF. The sampling efficiency was further improved to allow for reconstructing in the wild scenes [30]. UNISURF [22] also unifies surface and volume rendering but with occupancy fields. Similar to them, our work also only requires multi-view images for 3D reconstruction.
2.3 Implicit Representation with Sparse Views
Several works specifically deal with implicit representations given sparse views as it often happens in real-world applications. A common way to alleviate the requirements of dense inputs is to learn a generalized network from multiple scenes. These methods often contain a learned image encoder to encode inputs and then fuse multi-view features to assist the volume rendering process [37, 2, 11, 25, 16]. However, they require additional datasets for training and have potential domain generalization issues.
@@ 2.3 Implicit Representation with Sparse Views
Figure 2: Method Overview. We exploit the coarse geometry in radiance field training to guide its augmentation with sparse inputs. (left) For a surface point $\mathbf{v}$, we aggregate the color observations from all inputs to fit a spherical harmonics expansion, thus the pseudo labels for all the augmented rays passing through $\mathbf{v}$ can be obtained through querying the SH, (right) when generating augmented rays, we check their visibility and exclude those cannot be actually observed.
Another direction of work avoids pre-training through novel supervision in different aspects. DS-NeRF [5] and StructNeRF [4] supervise NeRF with sparse depth from structure-from-motion. The latter further exploits multi-view image and plane constraints for indoor scenes. DietNeRF [9] enforces semantic consistency of rendered patches with CLIP [24], which only provides inadequate high-level information. RegNeRF [14] regularizes the geometry and appearance of patches rendered from unobserved viewpoints, reducing the floating artifacts normally presented for limited input views. InfoNeRF [14] proposes regularizers for NeRF instead from the ray entropy perspective and can be flexibly plugged into other NeRF models. RayAug also has no pre-training procedure, but we differ from those works in that instead of introducing regularization or across scenes prior, we try to fit a radiance distribution of each surface point and use this as the augmented information. More specifically, we assume that the radiance distribution of a single point is a harmonic function and can be decomposed into a spherical harmonic expansion. Following this idea, we design a physically grounded pipeline to create precise pseudo labels for radiance supervision.
3 Preliminaries: NeuS
@@
NeuS [32] represents a 3D scene as two continuous functions both parametrized by a multi-layer perception (MLP). The first one is a signed-distance function that takes as input 3D position $\mathbf{x}=(x,y,z)$ and predicts its signed distance: $f_{1}(\mathbf{x})\rightarrow s$, and the other takes both $\mathbf{x}$ and its viewing direction $\mathbf{d}$ and predicts its view-dependent color: $f_{2}(\mathbf{x},\mathbf{d})\rightarrow\mathbf{c}$.
@@
Similar to NeRF [19], the color of a pixel in NeuS [32] only depends on the radiance along a ray with no other lighting factors. For a camera ray $\mathbf{r}(t)=\mathbf{o}+t\mathbf{d}$ with center $\mathbf{o}$ and direction $\mathbf{d}$, its color can be derived with volume rendering [12]. In practice, NeuS samples $N$ points along $\mathbf{r}(t)$ with $t=t_{1},...t_{N},t_{i}<t_{i+1}$ and approximate the per-pixel color with the following:
$\hat{\mathbf{C}}(\mathbf{r})$ $=\sum_{i=1}^{N}T_{i}\alpha_{i}\mathbf{c}_{i},$ (1)
$T_{i}$ $=\prod_{j=1}^{i-1}(1-\alpha_{j}),$ (2)
@@
$\mathcal{L}_{\mathrm{MSE}}=\sum_{\mathbf{r}}\|\hat{\mathbf{C}}(\mathbf{r})-\mathbf{C}(\mathbf{r})\|_{2}^{2}.$ (3)
@@
4 Methods
When limited images and corresponding training rays are given, from 3, only a small proportion of the 3D space has supervision, while the results of the rest space purely rely on the interpolation ability of the network. Therefore, the main goal of our paper is to augment the supervision signal of Implicit Neural Representation (INR) by inferring rays from unseen viewpoints based on limited inputs and the reconstructed INR in an accurate way. Particularly, we found that model the problem of inferring unseen radiance as fitting the radiance maps for spatial points, and the radiance map of each point could be estimated from the known radiance of that point based on our prior radiance map function.
In the following of this section, we introduce the details of our method. Our radiance augmentation algorithm is composed of three steps: Surface Ray Casting in 4.1, Visibility Checking in 4.2, and Radiance Map Estimation in 4.3.
4.1 Surface Ray Casting
Same to NeRF, the input to our method is a set of posed images $\{I_{i}\}_{i=1}^{N}$, where $N$ is the number of images. The original training set is denoted by =$\mathcal{R}=\{(\mathbf{r},\mathbf{c})\}$, in which $\mathbf{r}$ and $\mathbf{c}$ are the ray directions and ray color. Since our method relies on the rough scene geometry, we first train a radiance field for a few thousand steps and construct the mesh with marching cubes [17]. The vertices of this rough mesh are regarded as points lying on the surface. Then starting from these surface points $\mathbf{v}$, we shoot a number of rays $\mathbf{r}_{rand}\in\mathcal{R}$ in random directions. We denote the ray casted as ~$\tilde{\mathcal{R}}_{\mathbf{v}}$. We ensure that the angle between new rays and the surface normal of $\mathbf{v}$ is less than $\frac{\pi}{2}$, so these candidates for augmented rays are confident.
4.2 Visibility Check

### v5

# HarmonicNeRF: Geometry-Informed Synthetic View Augmentation for 3D Scene Reconstruction in Driving Scenarios

@@
Abstract.
In the realm of autonomous driving, achieving precise 3D reconstruction of the driving environment is critical for ensuring safety and effective navigation. Neural Radiance Fields (NeRF) have shown promise in creating highly detailed and accurate models of complex environments. However, the application of NeRF in autonomous driving scenarios encounters several challenges, primarily due to the sparsity of viewpoints inherent in camera trajectories and the constraints on data collection in unbounded outdoor scenes, which typically occur along predetermined paths. This limitation not only reduces the available scene information but also poses significant challenges for NeRF training, as the sparse and path-distributed observational data leads to under-representation of the scene’s geometry. In this paper, we introduce HarmonicNeRF, a novel approach for outdoor self-supervised monocular scene reconstruction. HarmonicNeRF capitalizes on the strengths of NeRF and enhances surface reconstruction accuracy by augmenting the input space with geometry-informed synthetic views. This is achieved through the application of spherical harmonics to generate novel radiance values, taking into careful consideration the color observations from the limited available real-world views. Additionally, our method incorporates proxy geometry to effectively manage occlusion, generating radiance pseudo-labels that circumvent the limitations of traditional image-warping techniques, which often fail in sparse data conditions typical of autonomous driving environments. Extensive experiments conducted on the KITTI, Argoverse, and NuScenes datasets demonstrate our approach establishes new benchmarks in synthesizing novel depth views and reconstructing scenes, significantly outperforming existing methods. Project page: https://github.com/Jiawei-Yao0812/HarmonicNeRF
Neural Radiance Fields, Sparse Views, Ray Augmentation, Autonomous Driving
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28-November 1, 2024, Melbourne, VIC, Australiadoi: 10.1145/3664647.3680851isbn: 979-8-4007-0686-8/24/10ccs: Computing methodologies Computer visionccs: Computing methodologies Computer graphics
1. Introduction
Figure 1. (a) illustrates the differing camera paths in both general surrounding scenes and specific driving scenarios. The camera trajectories in driving scenes are depicted as more linear and path-constrained, indicative of the typical movement patterns in autonomous driving data collection, as opposed to the more varied viewpoints found in general scenes. (b) shows our method’s effectiveness in dealing with the challenges of sparse and dynamic driving environments. Our reconstruction in (c) demonstrates significantly clearer and more accurate geometries
In recent years, with the rapid advancement of autonomous driving technology, 3D reconstruction has become a crucial component for ensuring precision in navigation and overall safety. Amidst this technological evolution, Neural Radiance Fields (NeRF) (Mildenhall et al., 2021) have emerged as a groundbreaking 3D reconstruction technique, gaining widespread attention for their ability to generate highly detailed and accurate models of complex environments. Unlike traditional Multi-View Stereo (MVS) (Yao et al., 2018; Chen et al., 2019; Furukawa et al., 2015) methods, NeRF utilizes deep learning (Cheng et al., 2024, 2023) to model environments, capable of synthesizing continuous and photorealistic images, thereby significantly enhancing reconstruction quality and visual fidelity. Furthermore, NeRF’s capability to synthesize intricate autonomous driving scenes (Hu et al., 2023; Li et al., 2023) enriches training datasets for autonomous systems, aiding in the improvement of their generalizability and decision-making quality.
Despite the remarkable performance of NeRF and its derivatives in targeted synthetic rendering datasets, their scene reconstruction performance significantly diminishes when applied to specific autonomous driving datasets such as KITTI (Geiger et al., 2012a) and NuScenes (Caesar et al., 2020). This decline is primarily attributed to the limited viewpoints provided by autonomous driving scenarios, failing to offer sufficient perspective information for high-quality synthesis. Additionally, the camera trajectories in autonomous driving are characterized by sparse viewpoints of single self-motion, differing from the object-centered forms previously encountered. The presence of moving objects in autonomous driving scenes introduces variations in lighting and geometry over time, resulting in artifacts in synthesized images and a decrease in reconstruction effectiveness. The original NeRF model does not account for moving objects, limiting its application in autonomous driving contexts.
To address the limitations of NeRF in modeling dynamic scenes undergoing motion, several scholars have proposed extensions to the NeRF network architecture. Innovations such as D-NeRF (Pumarola et al., 2021) and NRNeRF (Tretschk et al., 2021) for scenes with non-rigid deformations, and time-variable dynamic radiance fields like NSFF (Li et al., 2021) , NeRFlow (Du et al., 2021), DynamicNeRF (Gao et al., 2021), and RoDynRF (Liu et al., 2023) have been introduced. These adaptations incorporate temporal dimensions or consider non-rigid transformations to accommodate scene dynamics. Moreover, concerns regarding the original NeRF model’s size and its direct application to autonomous driving scenarios leading to significant artifacts and reduced visual fidelity have led to the development of solutions like Mip-NeRF (Barron et al., 2021a), which considers positions as conical sections of a light ray rather than points to mitigate aliasing effects on NeRF performance. NeRF in the Wild (Martin-Brualla et al., 2021) addresses varying environmental conditions by encoding appearances.
Although these methods have somewhat enhanced NeRF’s capability in dynamic scene reconstruction, they generally fail to fully address the challenges posed by the sparse viewpoints and path-distributed data collection inherent in autonomous driving scenes. In response to these issues, this paper introduces a novel geometry-guided ray augmentation technique specifically designed for sparse view scene reconstruction in driving scenarios. Our approach not only effectively manages challenges presented by dynamic objects and sparse viewpoints but also significantly improves scene reconstruction accuracy from sparse views through the innovative application of proxy geometry and spherical harmonics. Experiments conducted on challenging autonomous driving datasets such as KITTI, Argoverse (Chang et al., 2019), and NuScenes validate the superiority of our method, establishing new benchmarks in 3D scene reconstruction for the domain of autonomous driving. In summary, our contributions are as follows:
  • We leverage spherical harmonics to comprehensively integrate all color observations at a 3D point. This approach enables the generation of pseudo-labels that align with the natural distribution of radiance, enhancing the consistency and accuracy of scene illumination and texture.
  • We introduce a novel use of proxy geometry to address occlusion challenges during the ray augmentation process. This strategy ensures that point radiance is unconfounded, facilitating the reconstruction of more accurate surfaces from sparse viewpoints, a common scenario in autonomous driving environments.
@@ 1. Introduction
  • Our method is designed as a versatile, plug-and-play solution, compatible with existing sparse implicit neural surface reconstruction techniques. It demonstrates exceptional performance in sparse view reconstruction tasks, achieving superior results in driving scenarios without requiring additional data or extensive pre-training.
2. Related Work
@@
Neural Surface Reconstruction. Recently, neural implicit representations have demonstrated superior effectiveness in various tasks such as 3D object (Park et al., 2019; Mescheder et al., 2019; Chen and Zhang, 2019) and scene representation (Mildenhall et al., 2021; Sitzmann et al., 2019b, a), novel view synthesis (Mildenhall et al., 2021; Barron et al., 2021b; Wang et al., 2022) and multi-view 3D reconstruction (Yariv et al., 2020; Niemeyer et al., 2020; Wang et al., 2021; Oechsle et al., 2021). Among them, NeRF and its variants (Mildenhall et al., 2021; Barron et al., 2021b) demonstrated state-of-the-art performance for novel view synthesis with differentiable volumetric rendering. However, they require dense images for input and it is hard to extract high-quality geometry because the predicted density field lacks sufficient constraints. Other works utilizing surface rendering (Niemeyer et al., 2020) produce more accurate modeling but require additional constraints such as ground truth masks or depth priors for supervision (Zhang et al., 2021; Yariv et al., 2020). Several later studies combine the advantages of surface and volume representation, generating satisfying surfaces with only image inputs (Wang et al., 2021; Yariv et al., 2021; Oechsle et al., 2021). For example, NeuS (Wang et al., 2021) combined volume rendering with signed distance functions to circumvent the problem of insufficient surface constraint in the original NeRF. The sampling efficiency was further improved to allow for reconstructing in the wild scenes (Sun et al., 2022). UNISURF (Oechsle et al., 2021) also unifies surface and volume rendering but with occupancy fields. Similar to them, our work also only requires multi-view images for 3D reconstruction.
Large-Scale Scene Reconstruction. NeRF’s static assumptions falter in the dynamic, expansive scenes of autonomous driving, where direct application incurs artifacts and fidelity loss due to lighting and occlusion changes. Tackling NeRF’s capacity limitations, methods like Mega-NeRF (Turki et al., 2022) and Block-NeRF (Tancik et al., 2022) divide scenes into smaller, manageable NeRF segments, improving training efficiency and computational feasibility. Mip-NeRF (Barron et al., 2021a) addresses aliasing by representing positions as conical sections rather than points, enhancing rendering quality. Approaches like Block-NeRF’s adaptation of NeRF in the Wild (Martin-Brualla et al., 2021) encode varying appearances for realistic lighting condition reconstructions. Recent extensions such as LocalRF (Meuleman et al., 2023), READ (Li et al., 2023), and S-NeRF (Xie et al., 2023) further NeRF’s applicability in autonomous driving, dealing with changing geometries and appearances characteristic of these environments.
Implicit Representation with Sparse Views. In the realm of autonomous driving, reconstructing environments from sparse views is a pressing challenge due to the sporadic nature of data acquisition on the road. Two primary strategies have emerged to adapt NeRF for sparse input: leveraging pre-trained CNNs to extract image features from multi-view inputs or using sparse point clouds from Structure-from-Motion (SfM) for additional supervision.
PixelNeRF (Yu et al., 2021) innovatively simulates continuous neural scenes from a few input images, employing pre-trained layers of convolutional neural networks (CNNs) and bilinear interpolation. This method enhances the feature extraction process for each sampling point, fully utilizing the characteristics of input images. It then conveys the extracted feature points, spatial locations, and viewing directions to the NeRF network, enabling the construction of continuous static scenes from sparse image sets (Yu et al., 2021). The General Radiance Field (GRF) (Trevithick and Yang, 2021) adopts a similar approach to PixelNeRF but differs in that it operates in a standardized canonical space, offering more generality and versatility, particularly under the varying conditions encountered in autonomous driving scenarios. Point-NeRF (Xu et al., 2022) combines the explicit representation of point clouds with the implicit NeRF technique. It harnesses the strengths of both 3D representation forms and adapts efficiently to the surfaces of scenes, a crucial feature for capturing the complexities of driving environments (Chen et al., 2021).
Other methods such as GeoNeRF (Johari et al., 2022) and DietNeRF (Jain et al., 2021) propose novel supervision techniques that enrich the input data for NeRF with semantic consistency or spatial geometry from unobserved viewpoints. These methods mitigate issues like floating artifacts that are common when dealing with sparse views but often require additional datasets for training, which may introduce domain generalization challenges. HarmonicNeRF bypasses the need for pre-training entirely, diverging from the aforementioned works. Instead of regularization or relying on scene priors, we focus on fitting the radiance distribution for each surface point, using it as augmented information. We posit that the radiance distribution at a point is a harmonic function, decomposable into a spherical harmonic expansion. This insight leads us to develop a physically grounded pipeline to generate precise pseudo labels for radiance supervision, tailor-made for the dynamic and variable conditions present in autonomous driving data collection (Kim et al., 2022; Deng et al., 2022).
@@
Figure 2. Method Overview. We exploit the coarse geometry in radiance field training to guide its augmentation with sparse inputs. (left) For a surface point $\mathbf{v}$, we aggregate the color observations from all inputs to fit a spherical harmonics expansion, thus the pseudo labels for all the augmented rays passing through $\mathbf{v}$ can be obtained through querying the SH, (right) when generating augmented rays, we check their visibility and exclude those cannot be actually observed.
3. Preliminaries: NeuS
@@
NeuS (Wang et al., 2021) represents a 3D scene as two continuous functions both parametrized by a multi-layer perception (MLP). The first one is a signed-distance function that takes as input 3D position $\mathbf{x}=(x,y,z)$ and predicts its signed distance: $f_{1}(\mathbf{x})\rightarrow s$, and the other takes both $\mathbf{x}$ and its viewing direction $\mathbf{d}$ and predicts its view-dependent color: $f_{2}(\mathbf{x},\mathbf{d})\rightarrow\mathbf{c}$.
@@
Similar to NeRF (Mildenhall et al., 2021), the color of a pixel in NeuS (Wang et al., 2021) only depends on the radiance along a ray with no other lighting factors. For a camera ray $\mathbf{r}(t)=\mathbf{o}+t\mathbf{d}$ with center $\mathbf{o}$ and direction $\mathbf{d}$, its color can be derived with volume rendering (Kajiya and Von Herzen, 1984). In practice, NeuS samples $N$ points along $\mathbf{r}(t)$ with $t=t_{1},...t_{N},t_{i}<t_{i+1}$ and approximate the per-pixel color with the following:
(1) $\hat{\mathbf{C}}(\mathbf{r})$ $=\sum_{i=1}^{N}T_{i}\alpha_{i}\mathbf{c}_{i},$
(2) $T_{i}$ $=\prod_{j=1}^{i-1}(1-\alpha_{j}),$
@@
(3) $\mathcal{L}_{\mathrm{MSE}}=\sum_{\mathbf{r}}\|\hat{\mathbf{C}}(\mathbf{r})-\mathbf{C}(\mathbf{r})\|_{2}^{2}.$
@@
4. Methods
When limited images and corresponding training rays are given, from Section 3, only a small proportion of the 3D space has supervision, while the results of the rest space purely rely on the interpolation ability of the network. Therefore, the main goal of our paper is to augment the supervision signal of Implicit Neural Representation (INR) by inferring rays from unseen viewpoints based on limited inputs and the reconstructed INR in an accurate way. Particularly, we found that model the problem of inferring unseen radiance as fitting the radiance maps for spatial points, and the radiance map of each point could be estimated from the known radiance of that point based on our prior radiance map function.
In this work, we introduce HarmonicNeRF: Enhancing Neural Radiance Fields with Visibility-Driven Spherical Harmonics, a novel approach that significantly advances the capabilities of neural radiance fields by incorporating spherical harmonics for accurate radiance map estimation and employing visibility checks alongside surface ray casting for improved scene reconstruction from sparse viewpoints.
In the following of this section, we introduce the details of our method. Our radiance augmentation algorithm is composed of three steps: Surface Ray Casting in Section 4.1, Visibility Checking in Section 4.2, and Radiance Map Estimation in Section 4.3.
4.1. Surface Ray Casting
Same to NeRF, the input to our method is a set of posed images $\{I_{i}\}_{i=1}^{N}$, where $N$ is the number of images. The original training set is denoted by =$\mathcal{R}=\{(\mathbf{r},\mathbf{c})\}$, in which $\mathbf{r}$ and $\mathbf{c}$ are the ray directions and ray color. Since our method relies on the rough scene geometry, we first train a radiance field for a few thousand steps and construct the mesh with marching cubes (Lorensen and Cline, 1987). The vertices of this rough mesh are regarded as points lying on the surface. Then starting from these surface points $\mathbf{v}$, we shoot a number of rays $\mathbf{r}_{rand}\in\mathcal{R}$ in random directions. We denote the ray casted as ~$\tilde{\mathcal{R}}_{\mathbf{v}}$. We ensure that the angle between new rays and the surface normal of $\mathbf{v}$ is less than $\frac{\pi}{2}$, so these candidates for augmented rays are confident.
4.2. Visibility Check